# Generalisation in Multi-task Offline Q-Learning

### v1

@@ Abstract
We study offline multitask reinforcement learning in settings where multiple tasks share a low-rank representation of their action-value functions. In this regime, a learner is provided with fixed datasets collected from several related tasks, without access to further online interaction, and seeks to exploit shared structure to improve statistical efficiency and generalization. We analyze a multitask variant of fitted Q-iteration that jointly learns a shared representation and task-specific value functions via Bellman error minimization on offline data. Under standard realizability and coverage assumptions commonly used in offline reinforcement learning, we establish finite-sample generalization guarantees for the learned value functions. Our analysis explicitly characterizes how pooling data across tasks improves estimation accuracy, yielding a $1/\sqrt{nT}$ dependence on the total number of samples across tasks, while retaining the usual dependence on the horizon and concentrability coefficients arising from distribution shift. In addition, we consider a downstream offline setting in which a new task shares the same underlying representation as the upstream tasks. We study how reusing the representation learned during the multitask phase affects value estimation for this new task, and show that it can reduce the effective complexity of downstream learning relative to learning from scratch. Together, our results clarify the role of shared representations in multitask offline Q-learning and provide theoretical insight into when and how multitask structure can improve generalization in model-free, value-based reinforcement learning.
@@ 1 Introduction
In many real-world reinforcement learning (RL) applications, collecting new experience through online interaction is costly, time-consuming, or unsafe. Offline reinforcement learning addresses this challenge by learning decision-making policies from fixed datasets collected by previously deployed behavior policies, without requiring further interaction with the environment. In practice, such datasets often arise from multiple related tasks, motivating the study of offline multitask reinforcement learning, where a learner seeks to exploit shared structure across tasks to improve generalization and sample efficiency [10, 15, 11].
A common assumption in multitask RL is that tasks share an underlying low-dimensional representation, even though their rewards or dynamics may differ. Learning such shared representations can substantially reduce statistical complexity, particularly when the number of samples per task is limited. Empirically, representation sharing has been shown to improve transfer and generalization across tasks, but providing theoretical guarantees in the offline setting remains challenging due to distribution shift, limited coverage, and recursive error propagation through Bellman updates.
Most existing theoretical analyses of multitask representation learning in RL focus on model-based approaches, where shared structure is recovered through maximum-likelihood estimation of transition models or latent dynamics. In contrast, many practical algorithms are model-free and value-based, relying on fitted Q-iteration and Bellman error minimization. Despite their widespread use, theoretical guarantees for such methods in multitask, offline settings are comparatively limited.
In this work, we study offline multitask reinforcement learning through the lens of value-based methods. We analyze a multitask variant of fitted Q-iteration that jointly learns a shared representation and task-specific value functions from offline data via Bellman error minimization. We adopt a low-rank realizability assumption on the optimal action-value functions, where tasks share a common (unknown) representation but have task-dependent linear decoders [11]. Within this framework, we provide finite-sample generalization guarantees that explicitly quantify how pooling data across $T$ tasks improves statistical efficiency, yielding error rates that scale as $O(1/\sqrt{nT})$ in the number of samples per task $n$ and the number of tasks $T$, up to standard horizon- and coverage-dependent factors.
Beyond the upstream multitask setting, we also consider a downstream offline scenario in which a new task shares the same underlying representation as the source tasks. We analyze how reusing the representation learned during multitask training affects value estimation for the new task, and show that it can reduce the effective complexity of downstream learning compared to learning the representation from scratch.
@@
  • We provide a finite-sample analysis of multitask fitted Q-iteration in the offline, model-free setting under a shared low-rank representation assumption.
@@ 2 Related Work
Offline or batch reinforcement learning studies policy learning from fixed datasets without further interaction with the environment. Classical approaches rely on fitted Q-iteration [8, 21] and subsequent deep extensions such as BCQ, CQL, and IQL [9, 14, 13]. Theoretical analyses of offline RL have focused on characterizing sample complexity under concentrability assumptions [19, 2, 12]. Our work extends this line of research to the multitask setting, where shared structure across tasks is exploited to improve statistical efficiency in the offline regime.
@@
Representation learning for reinforcement learning has been studied as a means to transfer information across related tasks [3, 23, 16]. Recent theoretical progress has formalized multitask representation learning as a low-rank factorization of value or transition functions [1, 7]. These works typically assume access to interactive data or known models. Our formulation builds directly on this multitask structure but focuses on offline data settings and on value-based (Q-function) estimation rather than model estimation.
@@
Most existing theoretical treatments of multitask offline RL are model-based. The closest to our work is Domingues et al. [6], who propose an offline multitask framework that estimates shared low-rank transition models via maximum likelihood estimation (MLE). Their analysis provides generalization guarantees for multitask model learning but does not extend to model-free Q-learning. Our approach replaces the MLE oracle with a Bellman-error minimization oracle, leading to a model-free formulation and analysis that directly connects to practical Q-learning methods.
@@
Model-free RL methods, such as Q-learning and actor–critic algorithms, are known to suffer from instability when trained from offline data [17]. Recent theoretical results for single-task offline Q-learning [12, 20] establish finite-sample guarantees via Bellman error concentration. Our work extends these ideas to the multitask setting and develops new generalization bounds for shared representation learning under mean-squared Bellman error minimization, thereby unifying multitask representation learning and offline value-based RL.
@@ 3 Preliminary
Episodic MDP. An episodic discrete-time Markov Decision Process (MDP), represented by $\mathcal{M}=(\mathcal{S},\mathcal{A},H,P,r)$, where $\mathcal{S}$ is the state space, $\mathcal{A}$ is a finite action space. $K$ is defined as the cardinality of $\mathcal{A}$, and $H$ is the finite episode length, $P=\{P_{h}\}_{h\in H}$ are the state transition probability distributions with $P_{h}:\mathcal{S\times A}\to\Delta(\mathcal{S})$, and $r=\{r_{h}\}_{h\in[H]}$ are the deterministic reward functions with $r_{h}:\mathcal{S}\times\mathcal{A}\to[0,1]$. Following (Jiang et al., 2017; Sun et al., 2019), we assume that the initial state $s_{1}$ is fixed for each episode. A policy $\pi$ is a collection of $H$ functions $\{\pi_{h}:\mathcal{S}\to\mathcal{A}\}_{h\in[H]}$ where $\pi_{h}(s)$ is the action that the agent takes at state $s$ and at the $h$-th step in the episode. Given a starting state $s_{h}$, $s_{h^{\prime}}\sim(P,\pi)$ denotes a state sampled by executing policy $\pi$ under the transition model $P$ for $h^{\prime}-h$ steps and $\mathbb{E}_{(s_{h},a_{h})\sim(P,\pi)}[\cdot]$ denotes the expectation over states $s_{h}\sim(P,\pi)$ and actions $a_{h}\sim\pi$. Moreover, for each $h\in[H]$, we define the value function under policy $\pi$ when starting from an arbitrary state $s_{h}=s$ at the $h$-th time step as
@@
We define the action-value function, also called the $Q$ function, for a given state action pair, $(s,a)$ as
@@
This form of learning arises from the fact that may agents would need to perform multiple tasks in the same environment, which is determined by a common state and action spaces $\mathcal{S\times A}$. Each task $t\in[T]$ (from a set of $T$ tasks) can be represented by an episodic discrete time MDP: $(\mathcal{S},\mathcal{A},H,P^{(t)},r^{(t)})$. We note that the MDPs for each task share the same state space, action space and episode lengths, while differing in the transition probabilities and rewards. A policy $\pi$ in a multitask episodic MDP can also be redefined as a set of policies $\pi=\{\pi^{(t)}\}_{t\in[T]}$, such that $\pi^{(t)}$ is itself a collection $\{\pi_{h}^{(t)}:\mathcal{S}\to\mathcal{A}\}_{h\in[H]}$. Based on this we can also now use task specific value function as follows:
@@
And similarly we also have a task specific Q function:
@@ 3.1 Multitask Offline Q-Learning with Shared Representations
In multitask offline Q-learning, the agent is provided with offline datasets collected from $T$ related source tasks, denoted $\{\mathcal{D}^{(t)}\}_{t\in[T]}$. Each task $t$ is associated with an episodic Markov decision process $\mathcal{M}^{(t)}=(\mathcal{S},\mathcal{A},H,P^{(*,t)},r^{(*,t)})$. All tasks share a common state–action space and horizon $H$, but may differ in their transition dynamics $P^{(*,t)}$ or reward functions $r^{(*,t)}$. We assume that the tasks share a latent representation of the state space: there exists a feature encoder $\phi^{*}:\mathcal{S\times A}\to\mathbb{R}^{d}$ such that the optimal action-value functions satisfy
$Q_{h}^{(*,t)}(s,a)=\langle\phi^{*}(s,a),\omega_{h}^{(*,t)}\rangle$
for some task-specific decoders: $\omega_{h}^{(*,t)}$. This shared structure allows the agent to transfer information across tasks through a common embedding $\phi^{*}$. We adopt this low-rank structure as it is popular and has been emperically justified in previous work ([4]). The primary theoretical benefit of using a linear decoder is its minimal contribution to the overall function class capacity, which is crucial because this capacity scales rapidly demonstrated in Theorem 1.
The learning algorithm has access only to offline datasets
$\mathcal{D}^{(t)}=\bigcup_{h\in[H]}\mathcal{D}^{(t)}_{h},\quad\mathcal{D}^{(t)}_{h}=\{(s_{h}^{(i,t)},a_{h}^{(i,t)},r_{h}^{(i,t)},s_{h+1}^{(i,t)})\}_{i=1}^{n}$
collected under fixed behavior policies $\pi_{b}^{(t)}$ for each task. Each dataset therefore consists of transitions (state and action) $s_{h+1}^{(i,t)},a_{h}^{(i,t)}\sim P^{(*,t)}_{h},\pi_{b}^{(t)}$ and observed rewards $r_{h}^{(i,t)}$. The goal of upstream multitask learning is to learn a shared representation $\hat{\phi}$ and task-specific heads $\{\hat{\psi}^{(t)}\}_{t=1}^{T}$ that minimize the expected Q-Bellman error across tasks using only the offline data. Formally, at each step $h$, we minimise the following Q-bellman error:
$\hat{\mathcal{L}}(\phi,\{\psi^{(t)}\}_{t\in[T]})=\frac{1}{Tn}\sum_{t=1}^{T}\sum_{i=1}^{n}\Big(\langle\phi(s_{h}^{(i,t)},a_{h}^{(i,t)}),w_{h}^{(t)}\rangle-r_{h}^{(i,t)}-\max_{a^{\prime}}\langle\phi(s_{h+1}^{(i,t)},a^{\prime}),w_{h+1}^{(t)}\rangle)\Big)^{2}$
@@ 3.2 Role of the Concentrability Coefficient
A central difficulty in offline reinforcement learning is distribution shift between the data-generating distribution and the distributions induced by Bellman backups. Since learning is performed using samples from a fixed behavior policy, errors measured under the behavior distribution may be amplified when propagated backward through the Bellman operator.
@@
which upper bounds the density ratio between Bellman-induced distributions and the offline data distribution.
@@
This multitask empirical-risk objective performs joint representation learning and value-function regression across all tasks, providing a shared encoder that generalizes to unseen downstream environments. To ensure bounded sample complexity, we adopt the coverage condition standard in offline RL, assuming that each behavior policy $\pi_{b}^{(t)}$ sufficiently covers the state–action distribution induced by optimal or target policies. Under this assumption, our analysis establishes finite-sample generalization guarantees for the learned representation $\hat{\phi}$ and task-specific Q-functions $\hat{Q}^{(t)}$, showing that shared representations learned via multitask Q-learning achieve comparable sample efficiency to multitask model-based approaches. In Section 4.2, we formalize this result and provide our main finite-sample bound, demonstrating that the proposed multitask Q-learning framework admits strong theoretical guarantees in the offline setting.
Algorithm 1 Multi-task Fitted Q-Iteration
@@
4 Upstream Offline Multitask Representation Learning
In this section, we introduce our offline Multitask Q-Learning designed for upstream offline multitask RL in a low-rank Q-learner setting and describe its theoretical properties.
@@ 4.1 Multitask Fitted Q-Iteration Algorithm
The details of the algorithm MQRL (Multitask Q-Learning for Representation Learning) are depicted in Algorithm 1. The agent receives all offline datasets $\{\mathcal{D}^{(t)}\}_{t\in[T]}$ and jointly estimates the shared encoder $\phi_{h}$ and the task-specific decoders $\{\psi_{h}^{(t)}\}_{t\in[T]}$ by minimizing the empirical Bellman error across all tasks at each stage $h$:
@@
This objective acts as a multitask empirical-risk minimization oracle, serving as the offline Q-learning analogue of the maximum-likelihood oracle used in prior multitask model-based work (Agarwal et al., 2020b; Domingues et al., 2021). In practice, the optimization in 1 can be efficiently approximated whenever the function classes $\Phi$ and $\Psi$ are parameterized by neural networks or other differentiable architectures, allowing gradient-based minimization. The shared encoder $\phi_{h}$ aggregates statistical information across all tasks, thereby promoting representation sharing, while each decoder $\psi_{h}^{(t)}$ specializes to its respective task to approximate its local value function. In Algorithm 1 we prefer $\Theta$ to represent the combination of $\phi$ and $\omega_{h}^{(t)}$.

### v2

@@ Abstract
We study guarantees of offline reinforcement learning in a multitask setting Yu et al. (2021), focusing on Multitask Fitted Q-Iteration (MFQI) Borsa et al. (2016) under a shared low-rank structure in action-value functions. Given fixed datasets from multiple tasks over shared environments, we analyze how MFQI jointly learns a common representation and task-specific value functions via Bellman error minimization. Under standard offline RL assumptions, we derive finite-sample guarantees showing that pooling data across tasks improves estimation rates, achieving a $\mathcal{O}(1/\sqrt{nT})$ dependence on the total number of samples. This shows that a simple MFQI algorithm is competitive with similar algorithms in terms of sample complexity. We further consider transfer to a new task that shares the same representation, and show that reusing the learned features reduces the effective complexity of downstream learning. These results clarify when shared representations improve generalization in offline, value-based reinforcement learning. The code has been made available in [github.com/MandaKausthubh/PAC-Analysis-of-MFQI].
@@ 1 Introduction
In most real-world reinforcement learning (RL) applications, collecting new experience through online interaction is costly, time-consuming, or unsafe. Offline reinforcement learning addresses this challenge by learning decision-making policies from fixed datasets collected by previously deployed behavior policies, without requiring further interaction with the environment. In practice, such datasets often arise from multiple related tasks, motivating the study of offline multitask reinforcement learning Yu et al. (2021), where a learner seeks to exploit shared structure across tasks to improve generalization and sample efficiency (He et al., 2025; Kumar et al., 2023; Ishfaq et al., 2024; Kumar et al., 2022).
A common assumption Ishfaq et al. (2024); Cheng et al. (2022b); Oprescu et al. (2024) in multitask RL is that tasks share an underlying low-dimensional representation, even though their rewards or dynamics may differ. Learning such shared representations can substantially reduce statistical complexity, particularly when the number of samples per task is limited. Empirically and theoretically, representation sharing has been shown to improve transfer and generalization across tasks Cheng et al. (2022b); Ishfaq et al. (2024); Agarwal et al. (2020a), but providing theoretical guarantees in the offline setting remains challenging due to distribution shift, limited coverage, and recursive error propagation through Bellman updates.
Most existing theoretical analyses of multitask representation learning in RL focus on model-based approaches Ishfaq et al. (2024), where shared structure is recovered through maximum-likelihood estimation of transition models or latent dynamics. In contrast, many practical algorithms are model-free and value-based Zhu et al. (2024), relying on fitted Q-iteration and Bellman error minimization. One such model-free algorithm is Multitask-Fitted Q-iterations Borsa et al. (2016).
We analyse the offline multitask RL algorithm Multitask Fitted Q-Iterations Borsa et al. (2016) from the point of view of sample complexity. In our theoretical analysis, we consider imposing the low-rank realisability assumption (Assumption 4.2), under which all tasks share a common representation with task-dependent linear decoders. We also provide a short empirical discussion on the non-realisability cases as well. Our theoretical analysis shows us the advantage gained due to the multitask datasets, showing the competitiveness of MFQIBorsa et al. (2016) with other similar algorithms Agarwal et al. (2020a); Ishfaq et al. (2024); Cheng et al. (2022b); Jin et al. (2021).
We further consider a downstream offline setting where a new task shares the representation learned from the source tasks. By fixing this representation and learning only the task-specific value function, we obtain a simpler estimation problem that improves sample efficiency compared to learning both components from scratch. Similar downstream offline settings have been studied in previous literature Taylor and Stone (2009).
@@
  • We provide a finite-sample analysis of Multitask Fitted Q-iteration (MFQI)Borsa et al. (2016) in the offline, model-free setting under a shared low-rank representation assumption.
@@ 2 Related Work
Offline or batch reinforcement learning studies policy learning from fixed datasets without further interaction with the environment. Classical approaches rely on fitted Q-iteration (Ernst et al., 2005; Riedmiller, 2005) and subsequent deep extensions such as BCQ, CQL, and IQL (Fujimoto et al., 2019; Kumar et al., 2020; Kostrikov et al., 2021). Theoretical analyses of offline RL have focused on characterizing sample complexity under concentrability assumptions (Munos and Szepesvári, 2008; Antos et al., 2008; Jin et al., 2021). Our work extends this line of research to the multitask setting, where shared structure across tasks is exploited to improve statistical efficiency in the offline regime.
@@
Representation learning for reinforcement learning has been studied as a means to transfer information across related tasks (Bengio et al., 2013; Taylor and Stone, 2009; Lazaric, 2012). Recent theoretical progress has formalized multitask representation learning as a low-rank factorization of value or transition functions (Agarwal et al., 2020b; Du et al., 2019). These works typically assume access to interactive data or known models. Our formulation builds directly on this multitask structure but focuses on offline data settings and on value-based (Q-function) estimation rather than model estimation.
@@
Most existing theoretical treatments of multitask offline RL are model-based. The closest to our work is Ishfaq et al. (2024); Domingues et al. (2021), who propose an offline multitask framework that estimates shared low-rank transition models via maximum likelihood estimation (MLE). Their analysis provides generalization guarantees for multitask model learning but does not extend to model-free Q-learning. Our approach replaces the MLE oracle with a Bellman-error minimization oracle, leading to a model-free formulation and analysis that directly connects to practical Q-learning methods.
@@
Model-free RL methods, such as Q-learning and actor–critic algorithms, are known to suffer from instability when trained from offline data (Levine et al., 2020). Recent theoretical results for single-task offline Q-learning (Jin et al., 2021; Rashidinejad et al., 2022) establish finite-sample guarantees via Bellman error concentration. Our work provides an analysis of the popular method: MFQI Borsa et al. (2016) to the multitask setting and develops new generalization bounds for shared representation learning under mean-squared Bellman error minimization, thereby unifying multitask representation learning and offline value-based RL.
@@ 3 Preliminary
Episodic MDP. An episodic Markov Decision Process (MDP) is defined as $\mathcal{M}=(\mathcal{S},\mathcal{A},H,P,r)$, where $\mathcal{S}$ denotes the state space and $\mathcal{A}$ is a finite action space with cardinality $K=|\mathcal{A}|$. The horizon $H$ specifies the episode length. The transition dynamics are given by a collection of conditional distributions $P=\{P_{h}\}_{h\in[H]}$, where each $P_{h}:\mathcal{S}\times\mathcal{A}\to\Delta(\mathcal{S})$. The reward functions are $r=\{r_{h}\}_{h\in[H]}$ with $r_{h}:\mathcal{S}\times\mathcal{A}\to[0,1]$. We assume a fixed initial state $s_{1}$, following standard formulations and notations in the literature Ishfaq et al. (2024); Agarwal et al. (2020a).
A deterministic policy $\pi=\{\pi_{h}\}_{h\in[H]}$ maps states to actions at each stage, where $\pi_{h}:\mathcal{S}\to\mathcal{A}$. For a given policy $\pi$, we use $(s_{h},a_{h})\sim(P,\pi)$ to denote the state–action pair generated at step $h$ by following $\pi$ under dynamics $P$, and expectations are taken with respect to this induced trajectory distribution.
The value function at step $h$ is defined as
@@
and the corresponding action-value Q-function is
@@
We consider a collection of $T$ related tasks indexed by $t\in[T]$. Each task is modeled as an episodic MDP $\mathcal{M}^{(t)}=(\mathcal{S},\mathcal{A},H,P^{(t)},r^{(t)})$. The tasks share the same state space, action space, and horizon, but may differ in their transition kernels $P^{(t)}$ and reward functions $r^{(t)}$.
A multitask policy is defined as a collection $\pi=\{\pi^{(t)}\}_{t\in[T]}$, where each $\pi^{(t)}=\{\pi_{h}^{(t)}\}_{h\in[H]}$ governs decisions for task $t$. The corresponding task-specific value functions are
@@
Similarly, the task-specific Q-functions are
@@ 3.1 Multitask Offline Q-Learning with Shared Representations
In multitask offline Q-learning, we are given fixed datasets collected from $T$ related tasks, where each task corresponds to an MDP that shares a common state–action space but may differ in rewards and/or dynamics. The goal is to exploit shared structure across tasks to improve sample efficiency and generalization. This framing is similar to other general offline RL framing that exists in the literature Levine et al. (2020); Fujimoto et al. (2019); Kumar et al. (2020); Sutton and Barto (2012).
We assume that the optimal action-value functions admit a shared representation: there exists a feature encoder $\phi:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}^{d}$ such that for each task $t$,
$Q^{*,t}(s,a)=\langle\phi(s,a),w^{*,t}\rangle,$ (1)
where $w^{*,t}$ are task-specific parameters.
Learning proceeds via a multitask extension of Fitted Q-Iteration (FQI) Ernst et al. (2005), called Multitask Fitted Q-Iterations (MFQI) Borsa et al. (2016). At each iteration $k$, for every transition $(s,a,r,s^{\prime})$ in the dataset for task $t$, we construct bootstrapped targets using the current Q-function estimate:
$y^{(k)}_{t}(s,a)=r_{t}(s,a)+\gamma\max_{a^{\prime}}Q^{(k)}_{t}(s^{\prime},a^{\prime}).$ (2)
Given these targets, the update step reduces to a supervised multitask regression problem. Specifically, we jointly learn the shared representation $\phi$ and task-specific parameters $\{w_{t}\}_{t=1}^{T}$ by minimizing the empirical squared error across all tasks:
$\frac{1}{nT}\sum_{t=1}^{T}\sum_{i=1}^{n}\left(\langle\phi(s^{(i,t)},a^{(i,t)}),w_{t}\rangle-y^{(k)}_{t}(s^{(i,t)},a^{(i,t)})\right)^{2}.$ (3)
While the original MFQI Borsa et al. (2016) doesn’t explicitly state any loss function and creates a generic framework, we consider the Multitask loss to be as stated in 3 as it is widely used and as it provides a well-defined empirical risk minimization problem. This choice aligns with standard fitted Q-iteration and facilitates theoretical analysis through squared Bellman error minimization. This objective corresponds to minimizing the empirical Bellman residual at each iteration, while enforcing a shared representation across tasks. Crucially, the learning procedure alternates between (i) computing Bellman targets using the current Q-functions and (ii) solving a structured multitask regression problem, rather than directly optimizing a single fixed Bellman error objective.
The shared encoder $\phi$ couples learning across tasks, enabling statistical gains through data pooling and improving generalization in the low-sample regime.
@@ 3.2 Role of the Concentrability Coefficient
A central difficulty in offline reinforcement learning is distribution shift between the data-generating distribution and the distributions induced by Bellman backups. Since learning is performed using samples from a fixed behavior policy, errors measured under the behavior distribution may be amplified when propagated backward through the Bellman operator. To tackle this, we utilise the popular concentrability coefficient Munos (2003); Munos and Szepesvári (2008).
@@
which upper bounds the density ratio between Bellman-induced distributions and the offline data distribution. We assume a coverage condition such that the behavior distribution dominates all policy-induced distributions.
@@
This multitask empirical-risk objective performs joint representation learning and value-function regression across all tasks, providing a shared encoder that generalizes to unseen downstream environments. To ensure bounded sample complexity, we adopt the coverage condition standard in offline RL, assuming that each behavior policy $\pi_{b}^{(t)}$ sufficiently covers the state–action distribution induced by optimal or target policies. Under this assumption, our analysis establishes finite-sample generalization guarantees for the learned representation $\hat{\phi}$ and task-specific Q-functions $\hat{Q}^{(t)}$, showing that shared representations learned via multitask Q-learning achieve comparable sample efficiency to multitask model-based approaches. In Section 4.2, we formalize this result and provide our main finite-sample bound, demonstrating that the proposed multitask Q-learning framework admits strong theoretical guarantees in the offline setting. Similar methods have been used to derive bounds on offline RL in the literature Chen and Jiang (2019).
Algorithm 1 Multi-task Fitted Q-Iteration Borsa et al. (2016)
@@
4 Upstream Learning in MFQI
In this section, we review the offline multitask Q-learning algorithm presented in MFQI Borsa et al. (2016). This work builds upon, focusing on its formulation in the low-rank setting, and summarizes the theoretical properties relevant to our analysis.
@@ 4.1 Multitask Fitted Q-Iteration Algorithm
The details of the algorithm MQRL (Multitask Q-Learning for Representation Learning) Borsa et al. (2016) are depicted in Algorithm 1. The agent receives all offline datasets $\{\mathcal{D}^{(t)}\}_{t\in[T]}$ and jointly estimates the shared encoder $\phi_{h}$ and the task-specific decoders $\{\psi_{h}^{(t)}\}_{t\in[T]}$ by minimizing the empirical Bellman error across all tasks at each stage $h$:
@@
This objective acts as a multitask empirical-risk minimization oracle, serving as the offline Q-learning analogue of the maximum-likelihood oracle used in prior multitask model-based work Agarwal et al. (2020a); Domingues et al. (2021). In practice, the optimization in 4 can be efficiently approximated whenever the function classes $\Phi$ and $\Psi$ are parameterized by neural networks or other differentiable architectures, allowing gradient-based minimization. The shared encoder $\phi_{h}$ aggregates statistical information across all tasks, thereby promoting representation sharing, while each decoder $\psi_{h}^{(t)}$ specializes to its respective task to approximate its local value function. In Algorithm 1 we prefer $\Theta$ to represent the combination of $\phi$ and $\omega_{h}^{(t)}$.